OpenStudy (mathstudent55):

Solve the equation: \(\sqrt{x - 1} = x - 7\) @johnyy

OpenStudy (mathstudent55):

First, square both sides: \((\sqrt{x - 1})^2 = (x - 7)^2\) \(x - 1 = x^2 - 14x + 49\) \(0= x^2 - 15x + 50\) \(0 = (x - 10)(x - 5)\)

OpenStudy (mathstudent55):

x - 10 = 0 or x - 5 = 0 x = 10 or x = 5

OpenStudy (mathstudent55):

Since the equation had to be squared, we must check all solutions in the original equation. Check x = 10: \(\sqrt{10 - 1} = 10 - 7\) \(\sqrt{9} = 3\) \(3 = 3\) A true statement, so x = 10 is a true solution.

OpenStudy (anonymous):